# Reinforcement Learning and Machine ethics: a systematic review

### v1

@@ Discussion
We here present and analyze the results of a systematic review of articles that use reinforcement learning for machine ethics or consider the problem of machine ethics in reinforcement learning. The increased volume of papers in machine ethics RL in recent years points to a promising area of research. The main challenges for the RL in machine ethics community going forward is to organise itself around a taxonomy that will help authors position their efforts against existing work.
@@
The secondary challenge is to avoid “anecdotal ethics”. RL in machine ethics most of the time uses bottom-up and sometimes hybrid approaches (in the sense of [75]) to achieve artificial moral agency, to avoid developing methods around the moral sensitivities and judgments of its developers. The secondary challenge is thus to establish a firmer connection with the fields of moral philosophy and moral psychology [16]. The systematic evaluation and bench-marking of the abilities of reinforcement learning ethical agents remains an open problem.

### v2

@@ 6 Summary
We presented and analysed the results of a systematic review of articles that use RL for machine ethics or consider the problem of machine ethics in RL. The increased volume of papers in machine ethics RL in recent years points to a promising area of research. The main challenge for the RL in machine ethics community going forward is to organise itself around a taxonomy that will help authors position their efforts against existing work.
@@
The secondary challenge is to avoid “anecdotal ethics”. RL in machine ethics most of the time uses bottom-up and sometimes hybrid approaches (in the sense of [3]) to achieve artificial moral agency, to avoid developing methods around the moral sensitivities and judgments of its developers. The secondary challenge is thus to establish a firmer connection with the fields of moral philosophy and moral psychology [94]. The systematic evaluation and benchmarking of the abilities of reinforcement learning ethical agents remains an open problem.
Declarations
The authors declare no conflict of interest.